# Towards an Accessible, Noninvasive Micronutrient Status Assessment Method: A Comprehensive Review of Existing Techniques

### v1

@@ 3.2.4 Commercial Products
Method Form Factor Targets Analytes Evaluation Control Results Notes Source
Fluorometry Desktop device Vitamin A (retinol and retinyl esters) Breast milk Human subjects HPLC Weak correlation ($R^{2}$=0.59, p<0.001), but mean difference was “not statistically different from zero” Of major concern was the ability for the breast milk to satisfy the vitamin A requirements of children [4]
Quantitative immunoassay Desktop device Iron (ferritin), other tests Finger prick blood sample Not published Not published Not published Commercially available [38]
Immuno- fluorescence Desktop device Vitamin D (cholecalciferol) Serum Human subjects Abbott Alinity i immunoassay $R^{2}$ of 0.89; SE of 0.16 at 10 ng/mL, 0.19 at 12 ng/mL, and 0.35 at 30 ng/mL Standard error lower than control; recommends additional testing for excess status; target is not the clinically-accepted biomarker for Vitamin D [7]
3.3 Electrochemistry-Based Methods
Electrochemical analysis is a method that has been applied in literature to quantify levels of micronutrients in biofluids such as saliva, sweat, tears, urine, and blood. Huang et al. [66] summarizes the basis of electrochemical sensors with respect to vitamins, but we see work applying these ideas to minerals as well. The concentration of vitamins in an electrolyte (water or fat/organic solution that allows for the transfer of electrons) can be quantified by measuring electrical properties at a working electrode. The most common measurement techniques are voltammetry and amperometry. Voltammetry applies a varying voltage to the electrolyte and measures the resulting current, while amperometry applies a constant voltage and measures the resulting current over time [61]. Below, we dive into novel methods that were evaluated in biofluids, roughly divided into voltammetry and amperometry. For more detailed discussions, we direct the reader to recent reviews focusing on electrochemistry-based methods (e.g. [66, 126, 89, 147]). A summary of electrochemistry-based methods reviewed herein is found in Tables 4, 5, and 6.
3.3.1 Voltammetry
Voltammetry has been a popular method for the assessment of micronutrients in biofluids. Revin and John [143] propose a novel electrode for the simultaneous measurement of B2 (riboflavin), B9 (folic acid), and C (ascorbic acid) vitamins. Peak currents for each vitamin were well separated at mixtures of various concentrations. For vitamin C in particular, the tested linear range of the sensor was insensitive to the lower limit of sufficiency and below (Table 17). Additionally, the targets for vitamins B2 and B9 differed from their clinical gold-standard. Selectivity analysis showed that linearity in each vitamin was maintained even in the presence of elevated concentrations of all other vitamins. No interference from other common physiological interferents was found. Two plasma samples were tested before and after spiking with vitamin standards, with good recovery. Another electrode was developed by Jothimuthu et al. [77], targeting zinc. Interestingly, a sample pH of 6 was optimal for zinc assessment, which is more acidic than most biofluids (e.g. blood has a pH of 7.35-7.45). The zinc content in both acetate buffer and spiked, HCl-diluted serum was evaluated. A square wave voltammetric sensor was developed for the simultaneous measurement of glutathione (GSH), nicotinamide adenine dinucleotide (NADH) and folic acid (vitamin B9) [141]. A single urine and serum sample was collected for evaluation and spiked with known amounts of the targets. The authors did not present serum results.
Kim et al. [84] designed a square wave anodic-stripping voltammetric sensor has been used to monitor zinc in sweat during physical activity. The sensor itself is wearable, printed on tattoo transfer paper. The assessment was conducted with standard zinc solutions in a buffer medium. On-body experiments demonstrated the sensor’s ability to assess zinc in cycling-induced sweat. A linear calibration equation was used to quantify zinc concentrations, which were found to be close to the physiological range of zinc in sweat. Again, no gold-standard assessment methods for individual zinc status were used for comparison. However, this is less of a concern since this work explicitly focuses on providing insights into zinc excretion through sweat rather than determining in-body status.
Gao et al. [51] targeted zinc as well as copper using voltammetry. A wearable electrochemical sensor was created to assess Zn, Cd, Pb, Cu, and Hg ions in sweat and urine. Uniquely, the sensor incorporates skin temperature measurement for calibration and to compensate for the influence of temperature on electrochemical signals. This was important, as peak current was shown to increase with temperature. The device was developed with spiked synthetic sweat samples at concentrations an order of magnitude lower than is observed in blood (Table 19). Calibration curves demonstrate a linear relationship between peak current and concentration for the ions, but quantitative metrics are not reported. The authors conducted human studies for on and off-body measurements with ICP-MS as a control. The measured and controlled concentrations for Zn and Cu were similar, but statistical analysis of the results was not conducted. In Stanković et al. [162], a novel “boron-doped diamond electrode” for vitamin B12 (as cyanocobalamin) quantification is studied. Interference analysis showed a 10% signal change in the presence of a “10-fold excess of vitamin B6” [162]. The sensor was evaluated in four spiked urine samples, diluted, and pH adjusted to 2. In this case, the electrode targeted cyanocobalamin, which is the synthetic form of vitamin B12.
Sempionatto et al. [150] assumes a glasses form factor to conduct electrochemical analysis of tears, assessing the concentration of glucose, alcohol, vitamins B2, B6, and C. Tears were induced with menthol sticks before being collected and analyzed by the glasses. The sensor itself uses square wave voltammetry (SWV) for vitamin measurement, demonstrated only as a proof of concept. After a baseline was acquired with the sensor, tears were induced and analysis was conducted every 30 minutes for 2 hours after taking a multivitamin. Peak potentials emerged for each vitamin and were verified with known concentrations of vitamins added to baseline tear samples. Because this was a proof of concept, no comparison to gold standards with blood or efforts to quantify in-body vitamin levels were made.
A microfluidic, graphine-oxide-based sensor chip has been applied for the quantification of ferritin in serum [52]. Notably, the sample must be pumped through the sensor, where cyclic voltammetry is performed continuously with an external potentiostat. An evaluation was conducted with spiked serum samples (which were not sensitive to deficiency) and compared to ELISA. The sensor overestimated concentrations <100 mcg/L by $\sim$10and underestimated the larger concentration by $\sim$4%.
Sun et al. [163] focuses on low-cost and reusability in vitamin C assessment. Their device uses cyclic voltammetry to determine whether the vitamin C content present in the sample is normal or deficient (<4.93 mg/L). This deficiency threshold is greater than what is reported by the clinical literature, exceeding even the limit for sufficiency (Table 17). Interestingly, their device is also self-powered, using vitamin C as a biofuel. In a trial for scurvy detection (n=22), the device correctly determined the 4 deficient individuals (ground truth by HPLC). The authors also demonstrated its potential to screen for patients exhibiting a medical condition.
On-body electrochemical sensing of vitamins B6 (pyridoxine), C (ascorbic acid), D3 (calciol), and E (alpha-Tocopherol), as well as 9 amino acids and several macros in sweat, is enabled by Wang et al. [180]. Of the biomarkers targeted for each vitamin, only the target vitamin C aligns with the clinical gold-standard biomarker. Sweat is passively sampled using iontophoresis in a watch form factor. Voltammetry is used to detect vitamins indirectly. Quantitative results are not reported, but concentrations of vitamins appear to linearly correlate with peak height current density. The authors note the flexibility of this approach to the measurement of numerous other biomarkers. Human trials were conducted with healthy volunteers and patients but only targeted amino acids. This exemplifies the tendency of research to ignore micronutrients.
Lokesh Kumar et al. [101] developed a manganese dioxide nanoparticle–bimetallic metal-organic framework composite to detect vitamin D3 in spiked human plasma. Voltammetry measurements were compared to a gold standard for vitamin D detection, HPLC with ultraviolet detection (HPLC-UV), and obtained similar values. Seker et al. [148] designed a touch-based sensor that simultaneously monitors zinc and ascorbic acid (vitamin C) levels after supplementation. The technique measures fingertip sweat and uses SWV for zinc detection and potentiometric measurement for ascorbic acid detection. Lastly, Shi et al. [157] assessed an NFC-powered sensor for riboflavin (B2) in sweat. Selectivity testing was conducted by the addition of common sweat molecules into the standard, which did not significantly influence results. Uniquely, a pH sensor was incorporated into the device to account or the influence of pH on measurements. Human trials involved subjecting participants to exercise (n=1) or heat stress (n=2) to induce sweat after supplementation. Sweat samples were analyzed by the device and compared to HPLC sweat measurements for the exercise trials and fluorescence spectroscopy for the heat-stress trials. The time-dynamics of the device results followed the general trend of the control analysis methods, exhibiting more variance compared to urine results.
Table 4: Electrochemistry-Based Methods Using Voltammetry
@@ 3.3.1 Voltammetry
Method Form Factor Targets Analytes Evaluation Control Results Notes Source
Voltammetry Benchtop Vitamins B2 (riboflavin), B9 (folic acid), and C (ascorbic acid) Plasma Human subjects Spiked known concentrations >99% recovery of spiked concentrations Tested in 7.2 pH buffer, which is slightly lower than pH of blood ($\sim$7.4); direct riboflavin is not the gold-standard biomarker for vitamin B2; folic acid is the synthetic form of folate; Linear range for vitamin C did not reach below the lower limit for sufficiency [143]
Anodic stripping voltammetry Benchtop Zinc Serum Human subjects Spiked known concentrations Peak current decreased with concentration, but were lower in magnitude than buffer A sample pH of 6 was necessary for optimal performance [77]
SWV Benchtop Vitamin B9 (folic acid), GSH, and NADH Urine and serum Human subjects (urine only) Spiked known concentrations Accurate recovery of spiked concentration Simultaneous determination of targets; unclear whether pre-existing urine composition biased recovery; claims serum evaluation but this is not presented; folic acid is the synthetic form of folate [141]
Square wave anodic stripping voltammetry Wearable tattoo Zinc Sweat Zinc stock solutions Known zinc solutions $R^{2}$ of 0.999 for measured current vs stock solutions, LoD of 0.05 mcg/mL Zinc content in actual sweat from single participant was close to physiological range [84]
Square wave anodic stripping voltammetry Wearable patch Zn, Cd, Pb, Cu, and Hg ions Sweat and urine Human subjects ICP-MS Similar to control, but provided no statistical analysis Included temperature sensor to account for the influence of skin temperature on peak current [51]
SWV Benchtop Vitamin B12 (cyanocobalamin) Urine Diluted from human subjects Spiked known concentrations 98-105% recovery Cyanocobalamin is the synthetic form of B12 [162]
SWV and chronoamperometry Eyeglasses Vitamins B2, B6, and C, alcohol, and glucose Tears Human subjects Breathalyzer BAC, commercial glucometer, vitamin supplementation Correlations of 0.852 with BAC,  0.7 with glucometer, distinct voltage peaks for each vitamin Glucose and alcohol was main focus, vitamin assessment included as a proof of concept; exact form of each vitamin is not known [150]
Cyclic voltammetry Benchtop Iron (ferritin) Serum Spiked from human subjects ELISA $R^{2}$ of 0.966 and lower linearity than control; tended to overestimate concentrations <100 mcg/L Range of spiked concentrations was not sensitive to deficiency; studied the impact of pH and interferent compounds [52]
Cyclic voltammetry PoC Vitamin C Serum Human subjects HPLC $R^{2}$ of 0.984 (p<0.001) and 100% accuracy in deficiency detection Targets scurvy (extreme deficiency) but the threshold was set above sufficiency [163]
Voltammetry Wearable patch Vitamins B6 (pyridoxine), C (ascorbic acid), D3 (calciol), E (alpha-Tocopherol), and other macronutrients and amino acids Sweat Not reported for vitamins Not reported for vitamins Observable linear relationship between vitamin concentration and peak height current density Only vitamin C aligns with clinical standard, quantitative results for vitamins were not reported [180]
Voltammetry Benchtop Vitamin D3 Plasma Spiked from human subjects Spiked known concentrations and HPLC-UV LoD of 1.9 ng/mL; RSD of 0.3-2.6% and recovery of 96-102% Exact form of D3 (gold-standard 25(OH)D3/calcidiol or calciol) not reported [101]
SWV and potentiometric measurement PoC Zinc and vitamin C (ascorbic acid) Fingertip sweat Human subjects Supplementation Both micros could be analyzed over time simultaneously No comparison to clinical assessment of status or statistical analysis of results; Vitamin C range far exceeded physiological concentrations in plasma [148]
Differential pulse voltammetry Wearable patch Vitamin B2 (riboflavin) Sweat Human subjects HPLC $R^{2}$ of 0.9783 with sweat HPLC and 0.87 with urine fluorescence spectroscopy Urine included as comparison to sweat status; incorporates pH sensor to control for influence of pH on measurements; direct riboflavin is not the gold-standard biomarker for Vitamin B2 status [157]
3.3.2 Amperometry
Vitamins B2, B9, C, and D, as well as the mineral iron, have been assessed in biofluids using amperometry [106, 151, 190, 145, 191, 122]. Maiyalagan et al. [106] confront a major limitation of glassy carbon electrode-based sensors for vitamin B9 (folic acid): interference from vitamin C. Their nanofiber-modified electrode successfully avoided this interference. In evaluation, two serum samples were collected and evaluated before and after spiking with 4.41 mcg/L of folic acid. The peak current increased accordingly, allowing for greater than 99% recovery. Vitamin C is targeted by Sempionatto et al. [151], who deployed amperometry and immobilized ascorbate oxidase in a tattoo-based form factor. Tests on human subjects focused on the temporal characteristics of the current response, finding peak response in sweat 90 minutes after supplementation and a return to baseline 180 minutes after, in line with the plasma response of vitamin C. Tears and saliva are noted as other possible biofluids, with tears yielding a similar temporal profile (albeit with different peak currents). The authors also experimented with providing different amounts of vitamin C through both supplements and orange juice, claiming the response from sweat samples to increase in line with increasing vitamin C content. Crucially, no statistical analysis was conducted on the results of the experiments, such as correlations between intake and measured current. There are also no comparisons made to clinical reference methods of vitamin C assessment, though this is stated as a subject for future work. A wearable, electrochemical device to measure vitamin C levels in sweat, urine, and blood was proposed by Zhao et al. [190]. The device was used in a study where 6 male participants, aged 20-30, were given vitamin C as emergen-C brand supplements. The sensor is wearable, but no on-body measurements were made. Instead, urine and induced sweat were collected three hours post-intake and measured with the device. Blood samples were collected from a single participant in a separate study, analyzed with the device, and compared to results from urine and sweat. This research does not compare device results to any gold standard for vitamin C assessment; instead, it relies on intake, which (as we will see in Section 4) is a poor equivalent for in-body status due to individual differences in micronutrient absorption. In addition, we note that the vitamin C supplement, emergen-C, contains several other nutrients that could influence the results.
One paper proposes simultaneous measurement of vitamins C and D from a single saliva sample [145]. Their sensor combines an electrocatalytic vitamin C (ascorbic acid) amperometric assay and competitive vitamin D (25(OH)D3) immunoassay. Vitamin D is the clear focus of the paper, the vitamin C sensor gets little to no attention aside from some analysis of potential cross-talk between the two sensors. The sensor was applied in a study that supplemented vitamins to 3 participants and used the device to analyze saliva samples at increasing time intervals from intake. No gold-standard assessment method was used to evaluate the sensors, although the authors advocate for this evaluation and the development of truly quantitative sensors in future work. Another flexible, electrochemical sweat biosensor for vitamin C used polyaniline film modified with phytic acid [191]. The biosensor was validated with synthesized vitamin C samples of known concentrations. Human subjects were given supplements and had their sweat collected 3 times over 90 minutes. The sensor detected a general increase in current from the sweat samples over time, with variation across subjects. Saliva tests were also conducted. Peak current occurred 60 minutes after supplementation, in line with results from [151]. Another group designed a finger-actuated wirelessly-charging wearable that measures vitamin C and levadopa (a central nervous system agent) levels from sweat [122]. The system has a microfluidic chip with a self-driven pump and anti-reflux valve, a flexible wireless circuit board, and an associated smartphone app. They ran a study with five healthy participants whose sweat was collected and measured after exercising and ingesting vitamin C tablets as well as fava beans [122]. The results were not compared to clinical results as a gold standard baseline.
Table 5: Electrochemistry-Based Methods Using Amperometry

### v2

@@ 3.2.3. Smartphone-based Quantitative Assays
Method Platform Targets Analytes Evaluation Control Results Notes Source
Quantitative immunoassay Smartphone-based Vitamin D (calcidiol) Serum Known sample solutions and human subjects ELISA Errors “at same order” as ELISA No statistical analysis of results (Lee et al., 2014)
Quantitative immunoassay Smartphone-based Vitamin B12 (cyanocobalamin) Capillary blood sample Human Subjects Immulite 2000 Immunoassay Correlation of 0.93 with control,  85% specificity and  60% sensitivity for deficiency detection Used synthetic form of B12; poor accuracy outside of deficiency range (Lee et al., 2016b)
Quantitative immunoassay Smartphone-based Iron (total) Capillary serum sample Human Subjects Multi-Well Plate Spectrophotometry $R^{2}$ of 0.98 with control, CoV of 10.5% Conducted specificity testing using interferent analytes; considered toxicity (Serhan et al., 2020)
Quantitative immunoassay Smartphone-based Iron (total) Capillary blood sample Human Subjects Laboratory developed test: spectrophotometry-based technique Correlation plot with slope of 1.09, R2 of 0.96, and a mean bias of 5.3% Improved on control LoD; manually diluted some samples to represent a low iron concentration (Serhan et al., 2021)
Quantitative immunoassay Smartphone-based or portable device (TIDBIT) Iron (ferritin) Capillary blood sample Human Subjects Immulite 2000 Immunoassay Correlation of 0.92 with control, sensitivity of 0.9 for deficiency detection (Srinivasan et al., 2018)
Quantitative immunoassay Portable device (TIDBIT) Vitamin D3 (calcidiol) Serum and whole blood Commercial standards; human subjects Known solutions; LC-MS/MS CoV with standards of 2.63% at 34 ng/mL and 11.2% at 0 ng/mL; $R^{2}$ of 0.91 for serum tests, 0.94 for capillary blood tests (Vemulapati et al., 2017)
Colorimetric assay Paper-based Iron (total) Urine Human subjects AAS RSD of 9.5% Urine citric acid was found to interfere with results (Ferreira et al., 2021)
Colorimetric assay Paper-based Iron (Fe3+) Serum Spiked, diluted from human subjects Spiked known concentration Error of 3.7% and RSD of 1% LoD of 0.3 $\mu$g/mLl (Dortez et al., 2023)
Colorimetric assay Paper-based Iron (Fe3+) Saliva Spiked pooled commercial saliva Spiked known concentration $R^{2}$ of 0.99 It is debated whether salivary iron reflects circulating status (Prakobdi et al., 2024)
Opto-electronic immunoassay Card Vitamin A (RBP), Iron (ferritin), and CRP Whole blood Human subjects ELISA CVs of 2.5% for ferritin, 10.8% for RBP, and 3.9% for CRP Cutoff for iron deficiency was insensitive to upper limit; RBP is not the clinically-accepted biomarker for vitamin A status; Device transmits results to a smartphone over NFC (Lee et al., 2016a)
Multiplexed quantitative assay Portable device (TIDBIT) Vitamin A (RBP), Iron (ferritin), and CRP Serum Human subjects ELISA $R^{2}$ of 0.56 for RBP, 0.92 for ferritin, 0.88 for CRP Ferritin range does not cover anemia; RBP is not the clinically-accepted biomarker for vitamin A status; RMSE for RBP was 21 $\mu$g/mL (Lu et al., 2017)
3.2.4. Commercial Products
Commercially-available devices have emerged in recent years to provide point of care (PoC) testing for some micronutrients (Abebe et al., 2019; Diagnostics, 2021; Fiddler et al., 2023; Albrecht et al., 2021). One study has explored the utility of a commercial iCheck FLUORO device to assess vitamin A concentrations in human milk (human milk vitamin A or HMVA) (Abebe et al., 2019). HMVA is critical since it is the primary source of vitamin A for breastfeeding children. If there is a vitamin A deficiency in human milk, it is likely to cause developmental issues for a child. The authors collected human milk samples and socio-demographic and anthropomorphic data from lactating mothers in the Mecha district, Ethiopia (n=104). This region was selected because prior studies applying this device for HMVA assessment recommended further investigation of populations at greater risk of vitamin A deficiency. Concentrations of vitamin A in human milk were measured by iCheck FLUORO and compared to HPLC. The commercial device was found to overestimate low HMVA concentrations and had a weak overall correlation with HPLC results. Therefore, the paper concluded that studies which assess vitamin A intake among breast-feeding children in developing countries should not assume average HMVA. It was argued that devices like the FLUORO are needed to monitor HMVA status, especially for intervention programs that typically assume average HMVA. Still, they must be “reliable across a range of HMVA concentrations” (Abebe et al., 2019).
Albrecht et al. (2021) likewise studied the efficacy of the Quidel Inc Sofia fluorescent immunoassay for serum vitamin D (as calciol). It should be noted that calciol is an inactive form of vitamin D, distinct from the optimal reference standard of calcidiol for vitamin D status. The assay was analyzed by the Sofia Analyzer, a PoC device for immunoassay analysis. A total of 324 samples were collected and 296 were used (229 female). Additionally, 433 tests were run using both frozen (208) and fresh samples (88). Notably, the researchers also assessed random error and inter-operator reliability for the device. Because only one sample had a concentration above 100 ng/mL, the authors recommend additional testing for concentrations above 80 ng/mL (well into the range of toxicity; Table LABEL:tab:_assessment).
Bloom Diagnostics is a home use ‘lab’ device that analyzes single-use qualitative test strips to quantitatively assess the status of in-vitro (in-body) biomarkers, similar to the TIDBIT (Vemulapati et al., 2017; Lu et al., 2017; Diagnostics, 2021). Tests available for Bloom include thyroid-stimulating hormone (TSH), ferritin, CRP, and estimated glomerular filtration rate with cystatin C. While Bloom approaches the goal of accessible nutrition assessment, its assays still require the user to collect a sample themselves. Depending on the test, this could involve a finger prick, coaxing the blood into a collection tube, and depositing it properly onto the assay. VitaScan is another commercial PoC device that tests for iron deficiency, and they validate results against the clinical optimal reference standard for in-body iron measurement (Fiddler et al., 2023). The device is not yet released, but it is planned to assess vitamins B12, D, and A and CRP in the future. The method is still invasive as it utilizes capillary blood obtained from a finger prick and also requires the user to obtain the sample themselves.
Table 4. Commercial Assay-Based Methods
@@ 3.2.4. Commercial Products
Method Platform Targets Analytes Evaluation Control Results Notes Source
Fluorometry Portable device Vitamin A (retinol and retinyl esters) Breast milk Human subjects HPLC Weak correlation ($R^{2}$=0.59, p¡0.001), but mean difference was “not statistically different from zero” Of major concern was the ability for the breast milk to satisfy the vitamin A requirements of children (Abebe et al., 2019)
Immuno- fluorescence Portable device Vitamin D (cholecalciferol) Serum Human subjects Abbott Alinity i immunoassay $R^{2}$ of 0.89; SE of 0.16 at 10 ng/mL, 0.19 at 12 ng/mL, and 0.35 at 30 ng/mL Standard error lower than control; recommends additional testing for excess status; target is not the clinically-accepted biomarker for Vitamin D (Albrecht et al., 2021)
Quantitative immunoassay Home use device Iron (ferritin), other tests Capillary blood sample Not published Not published Not published Commercially available (Diagnostics, 2021)
3.2.5. Comparison
Although they are the most accessible and noninvasive, sweat-based assays are generally not a viable alternative to clinical methods of micronutrient status assessment given the current clinical knowledge about the correlation between sweat and blood concentrations of micronutrients. A critical tradeoff exists between accessibility and accuracy, yet until the clinical literature has reached an agreement on which alternative biosamples (i.e. not blood or urine) are most appropriate for the assessment of in-body status, it may be more productive to pursue accessibility through device design rather than through the biosample analyzed so as to not sacrifice clinical-relevance.
Microfluidics and smartphone-based quantification systems may deliver on accessibility and noninvasiveness. The former enables analysis on a small volume of a biosample, meaning that less of the sample needs to be collected (via potentially invasive means in the case of blood). The latter decreases reliance on commercial assay analyzers, which may only operate on proprietary assays, and enables accessibility through portability and cost-effectiveness.
As opposed to other types of assays (e.g. immunoassays), colorimetric assays have been overwhelmingly applied to the assessment of iron (with one expanding to sweat vitamin C, calcium, and zinc). This may indicate the poor versatility of colorimetric assays, whereas other assay techniques are able to assess a wider variety of biomarkers, largely through adjustments to the antibodies used in the assay. The trend of multiplexed assays in both sweat-based and non-sweat-based sensors is therefore a promising means to balance the pros and cons of different assay types. While multiplexed solutions may increase cost, we argue that the potential to assess multiple biomarkers at once significantly decreases the burden on the patient and the clinician caused by the need to use multiple assays/devices.
3.3. Electrochemistry-Based Methods
Electrochemical analysis is a method that has been applied in literature to quantify levels of micronutrients in biofluids such as saliva, sweat, tears, urine, and blood. Huang et al. (2021) summarizes the basis of electrochemical sensors with respect to vitamins, but we see work applying these ideas to minerals as well. The concentration of vitamins in an electrolyte (water or fat/organic solution that allows for the transfer of electrons) can be quantified by measuring electrical properties at a working electrode. The most common measurement techniques are voltammetry and amperometry. Voltammetry applies a varying voltage to the electrolyte and measures the resulting current, while amperometry applies a constant voltage and measures the resulting current over time (Harvey, 2024). Below, we dive into novel methods that were evaluated in biofluids, roughly divided into voltammetry and amperometry. For more detailed discussions, we direct the reader to recent reviews focusing on electrochemistry-based methods (e.g. (Huang et al., 2021; Pakeeza et al., 2024; Kirazoğlu and Benli, 2023; Sardar et al., 2023)). A summary of electrochemistry-based methods reviewed herein is found in Tables 5, 6, and 7.
3.3.1. Voltammetry
Voltammetry has been a popular method for the assessment of micronutrients in biofluids. Revin and John (2012) proposed a novel electrode for the simultaneous measurement of vitamins B2 (riboflavin), B9 (folic acid), and C (ascorbic acid). Peak currents for each vitamin were well separated at mixtures of various concentrations. For vitamin C in particular, the tested linear range of the sensor was insensitive to the lower limit of sufficiency and below (Table LABEL:tab:_assessment). Additionally, the analyzed biomarkers for vitamins B2 and B9 differed from their clinical optimal reference standard. Selectivity analysis showed that linearity in each vitamin was maintained even in the presence of elevated concentrations of all other vitamins. No interference from other common physiological interferents was found. Two plasma samples from a clinical laboratory were diluted and tested before and after spiking with vitamin standards, with good recovery. Another electrode was developed by Jothimuthu et al. (2013), examining zinc. Interestingly, a sample pH of 6 was optimal for zinc assessment, which is more acidic than most biofluids (e.g. blood has a pH of 7.35-7.45). The zinc content in both acetate buffer and spiked, HCl-diluted serum was evaluated. A square wave voltammetric sensor was developed for the simultaneous measurement of glutathione (GSH), nicotinamide adenine dinucleotide (NADH) and folic acid (vitamin B9) (Raoof et al., 2015). A single urine and serum sample was collected for evaluation and spiked with known amounts of the targets. The authors did not present serum results.
Kim et al. (2015a) designed a square wave anodic-stripping voltammetric sensor to monitor zinc in sweat during physical activity. The sensor itself was wearable, printed on tattoo transfer paper. The assessment was conducted with standard zinc solutions in a buffer medium. On-body experiments (7 participants, 5 male) demonstrated the sensor’s ability to assess zinc in cycling-induced sweat, which was found to be close to the physiological range. Again, no clinical reference standard assessment methods for individual zinc status were used for comparison. However, this is less of a concern since this work explicitly focuses on providing insights into zinc excretion through sweat rather than determining in-body status.
Gao et al. examined zinc as well as copper using voltammetry Gao et al. (2016). A wearable electrochemical sensor was created to assess Zn, Cd, Pb, Cu, and Hg ions in sweat and urine. Uniquely, the sensor incorporated skin temperature measurement for calibration, and to compensate for the influence of temperature on electrochemical signals. This was important, as peak current was shown to increase with temperature. The device was developed with spiked synthetic sweat samples at concentrations an order of magnitude lower than is observed in blood (Table LABEL:tab:_assessment). Calibration curves demonstrate a linear relationship between peak current and concentration for the ions, but quantitative metrics were not reported. The authors conducted a human study with a single participant for on and off-body measurements with ICP-MS as a control. The measured and controlled concentrations for Zn and Cu were similar, but statistical analysis of the results was not conducted. In Stanković et al. (2016), a novel “boron-doped diamond electrode” for vitamin B12 (as cyanocobalamin) quantification was studied. Interference analysis showed a 10% signal change in the presence of a “10-fold excess of vitamin B6” (Stanković et al., 2016). The sensor was evaluated in four spiked urine samples, diluted, and pH adjusted to 2. In this case, the electrode analyzed cyanocobalamin, which is the synthetic form of vitamin B12.
Sempionatto et al. (2019) developed an eyeglasses-based platform to conduct electrochemical analysis of tears, assessing the concentration of glucose, alcohol, vitamins B2, B6, and C. Tears were induced with menthol sticks before being collected and analyzed by the glasses. The sensor itself used square wave voltammetry (SWV) for vitamin measurement, demonstrated only as a proof of concept. After a baseline was acquired with the sensor, tears were induced from 3 participants and analysis was conducted every 30 minutes for 2 hours after taking a multivitamin. Peak potentials emerged for each vitamin and were verified with known concentrations of vitamins added to baseline tear samples. Because this was a proof of concept, no comparison to optimal reference standards with blood or efforts to quantify in-body vitamin levels were made.
A microfluidic, graphine-oxide-based sensor chip has been applied for the quantification of ferritin in serum (Garg et al., 2020). Notably, the sample must be pumped through the sensor, where cyclic voltammetry was performed continuously with an external potentiostat. An evaluation was conducted with spiked serum samples (which were not sensitive to deficiency) and compared to ELISA. The sensor overestimated concentrations ¡100 $\mu$g/L by $\sim$10and underestimated the larger concentration by $\sim$4%.
Sun et al. (2021) focused on reusability in vitamin C assessment. Their device used cyclic voltammetry to determine whether the vitamin C content present in the sample is normal or deficient (¡4.93 mg/L). This deficiency threshold was greater than what is reported by the clinical literature, exceeding even the limit for sufficiency (Table LABEL:tab:_assessment). Interestingly, their device was also self-powered, using vitamin C as a biofuel. In a trial for scurvy detection (n=22), the device correctly determined the 4 deficient individuals (ground truth by HPLC). The authors also demonstrated its potential to screen for patients exhibiting a medical condition, identifying 30 patients (total sample size unclear) who suffered from vitamin C deficiency during routine checkups. Of these, ¿85% had a medical condition associated with inflammation and oxidative stress.
On-body electrochemical sensing of vitamins B6 (pyridoxine), C (ascorbic acid), D3 (calciol), and E (alpha-Tocopherol), as well as 9 amino acids and several macros in sweat, was enabled by Wang et al. (2022). Of the biomarkers examined for each vitamin, only the vitamin C biomarker aligned with the clinical optimal reference standard biomarker. Sweat was passively sampled using iontophoresis in a watch-based platform. Voltammetry was used to detect vitamins indirectly. Quantitative results were not reported, but concentrations of vitamins appeared to linearly correlate with peak height current density. The authors noted the flexibility of this approach to measure numerous other biomarkers. Human trials were conducted with healthy volunteers and patients but only examined amino acids. This exemplifies the tendency of mainstream research to ignore micronutrients.
Lokesh Kumar et al. (2023) developed a manganese dioxide nanoparticle–bimetallic metal-organic framework composite to detect vitamin D3 in spiked human plasma. Voltammetry measurements were compared to a optimal reference standard for vitamin D detection, HPLC with ultraviolet detection (HPLC-UV), and obtained similar values. Seker et al. (2024) designed a touch-based sensor that simultaneously monitored zinc and ascorbic acid (vitamin C) levels after supplementation. The technique measured fingertip sweat and uses SWV for zinc detection and potentiometric measurement for ascorbic acid detection. Lastly, Shi et al. (2024) assessed an NFC-powered sensor for riboflavin (B2) in sweat. Selectivity testing was conducted by the addition of common sweat molecules into the standard, which did not significantly influence results. Uniquely, a pH sensor was incorporated into the device to account or the influence of pH on measurements. Human trials involved subjecting participants to exercise (n=1) or heat stress (n=2) to induce sweat after supplementation. Sweat samples were analyzed by the device and compared to HPLC sweat measurements for the exercise trials and fluorescence spectroscopy for the heat-stress trials. The time-dynamics of the device results followed the general trend of the control analysis methods, exhibiting more variance compared to urine results.
Table 5. Electrochemistry-Based Methods Using Voltammetry